OpenStudy (anonymous):

Can you help me with this integral by parts and explain me x^3cos^4(x/8) dx

OpenStudy (castiel):

Are you sure that's the correct integral? Because this one looks really nasty

OpenStudy (anonymous):

There's a variety of ways to approach this. I'd suggest using the half-angle formula to reduce the power of the cosine. \[\begin{align*}\cos^4\frac{x}{8}&=\left(\frac{1+\cos\dfrac{2x}{8}}{2}\right)^2\\\\ &=\frac{1}{4}+\frac{1}{2}\cos\frac{x}{4}+\frac{1}{4}\cos^2\frac{x}{4}\\\\ &=\frac{1}{4}+\frac{1}{2}\cos\frac{x}{4}+\frac{1}{4}\left(\frac{1+\cos\dfrac{2x}{4}}{2}\right)\\\\ &=\frac{3}{8}+\frac{1}{2}\cos\frac{x}{4}+\frac{1}{8}\cos\frac{x}{2} \end{align*}\] Integral: \[\begin{align*} \int x^3\cos^4\frac{x}{8}~dx&=\frac{3}{8}\int x^3~dx+\frac{1}{2}\int x^3\cos\frac{x}{4}~dx+\frac{1}{8}\int x^3\cos\frac{x}{2}~dx\\\\ &=\frac{3}{32}x^4+\frac{1}{2}\int x^3\cos\frac{x}{4}~dx+\frac{1}{8}\int x^3\cos\frac{x}{2}~dx \end{align*}\] For the cosine integrals, when you integrate by parts be sure to use \[\begin{matrix} u=x^3&&&dv=\cos\frac{x}{n}\\\\ du=3x^2~dx&&&v=n\sin\frac{x}{n} \end{matrix}\] where \(u\) will be the power of \(x\) and \(v\) will be the trig factor.